OpenStudy (anonymous):

Admission to the museum is $12 for children and $15 for adults. On a certain day, 2450 people go to the museum and $33,456 is collected. How many children and how many adults went to the museum? ____children ____adults

OpenStudy (anonymous):

@Directrix

Directrix (directrix):

Assignment of variables: x represents children y represents adults

OpenStudy (anonymous):

i might know it

Directrix (directrix):

x + y = ?

OpenStudy (anonymous):

can i tell you what i think?

OpenStudy (anonymous):

nvm

Directrix (directrix):

I don't know the answer so you can tell me but we still will need to work this problem. I have worked this problem at least 10 times before but do not have the answers in memory.

Directrix (directrix):

x + y = ?

OpenStudy (anonymous):

ok so x+y=2450

Directrix (directrix):

Each child ticket is worth how much money?

OpenStudy (anonymous):

$12 for children

OpenStudy (anonymous):

$15 for adults

OpenStudy (anonymous):

btw my computer is acting dumb so.....yea

Directrix (directrix):

Child revenue is 12x Adult revenue is ?

OpenStudy (anonymous):

15y

Directrix (directrix):

The sum of child revenue and adult revenue is what?

OpenStudy (anonymous):

you mean 12+15?

Directrix (directrix):

>>btw my computer is acting dumb so.....yea So is mine but it's not the computer. It is this site.

OpenStudy (anonymous):

oh ok lol

OpenStudy (anonymous):

do you want me to add 12+15?

Directrix (directrix):

No.

OpenStudy (anonymous):

multply them? divde them?

Directrix (directrix):

So, the sum of these two is what?

Directrix (directrix):

Child revenue is 12x Adult revenue is 15y This is algebra. We use variables. Algebra is the science of finding x. 12x + 15y = ?

Directrix (directrix):

Look at the problem. How much money was taken in that day at the museum?

OpenStudy (anonymous):

$33,546

OpenStudy (anonymous):

$33,456

Directrix (directrix):

x + y = 2450 12x + 15y = 33456 Take it to the wolf for solution.

OpenStudy (anonymous):

ok

Directrix (directrix):

x represents children y represents adults so don't put the children number in the adult slot and vice versa

Directrix (directrix):

Did you lose the wolf? http://www.wolframalpha.com/

OpenStudy (anonymous):

i got something weird

OpenStudy (anonymous):

i don't get it

OpenStudy (anonymous):

1098 = Children 1352 = Adults

Directrix (directrix):

Post the two equations you entered in the same way you entered them.

OpenStudy (anonymous):

who is that?

Directrix (directrix):

1098 = Children 1352 = Adults is what I got. Did you use the Wolf or cherry-pick from the options?

OpenStudy (anonymous):

what is cherry pick?

Directrix (directrix):

____children ____adults Don't put the children number in the adult slot. Focus.

OpenStudy (anonymous):

can you send me the link plz

Directrix (directrix):

>>what is cherry pick? Google it up. Time to increase vocab.

Directrix (directrix):

>>can you send me the link plz link for what?

OpenStudy (anonymous):

i used wolf

Directrix (directrix):

Close, post, and move on.